INAF LBC+MODS queue observing, 2011 October 21-22

Observer: K. Boutsia, M. Fumana, P. Santini
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LBC + MODS

Summary

Thin clouds all over the sky in the afternoon. We kept having clouds passing through for the rest of the night. Some OBs need to be repeated because of this, thus ~1hr was lost to weather.
30min were lost because while switching to MODS the secondary ripped and we had to call for assistance resetting it.
After we recentered the mask and started MODS spectroscopy again, the LDG rotator faulted and we lost another 30min (net time for the rotator - not considering the time it took to recenter the mask).
In the dwan we took calibration data with MODS (flats +arcs) for both masks used by the Donnarumma program (b1 and b2).

Data Summary

Project Objects Filter ExpTime Notes
Bragaglia NGC1817 Bbess - Vbess / Ibess 20min the seeing was <1'' but sky not photometric
Clementini AndXXV Bbess - Vbess 2hr 20min 1 OB out of specs because of clouds
Clementini AndXXVII Bbess - Vbess 2hr 20min 1 OB out of specs because of clouds
STD Stetson_L92 Bbess - Vbess / Ibess   1 position and 4 positions - airmass 1.2
STD SA113 Bbess - Vbess / Ibess   4 positions - airmass 1.2
Donnarruma MAC0717b1 r_sdss 1hr15min there are some clouds - seeing is <1.2''

Details

12:52 Test SkyFlats
  • target moved away from the Galactic plane (16:00:00,+50:30:00)
  • the new position looks better (not as many stars)
  • BV 0 (1.5sec-1sec) - scale 0.5 - B 50000 counts / V 30000 counts (01:05)
  • BV 0 (1.5sec-1sec) - scale 0.5 - B 30000 counts / V 20000 counts (01:07)
  • BV 0 (1.5sec-1sec) - scale 0.5 - B 18000 counts / V 13000 counts (01:10)
  • tomorrow we can try BV (1sec-1sec) - but B drops quickly than V
  • BV 180 (1.5sec-1sec) scale 1 - B 11000 counts / V 8000 counts (01:12)
  • BV 180 (1.5sec-1sec) scale 1 - B 10000 counts / V 9000 counts (01:13)
  • BV 180 (1.5sec-1sec) scale 1.5 - B 10000 counts / V 9000 counts (01:15)
  • BI 180 (3sec-1sec) scale 3.5 - B 21000 counts / I 40000 counts (01:21)
  • BI 180 (3sec-1sec) scale 3.5 - B 17000 counts / I 35000 counts (01:22)
  • BI 0 (3sec-1sec) scale 4 - B 8000 counts / I 17000 counts (01:27)

01:30 Slew to pointing star WT10_219
  • dofpia,/first
  • dofpia (01:41)

01:46 lbcrangebal on WT10_219

01:50 Slew to SA113 for standard calibration (elev~53 , airmass=1.2)
  • dither in 4 positions

01:56 Slew to M31_AndXXV_400s_12 * dofpia

02:02 M31_AndXXV_400s_12
  • seeing measured on first images ~0.88 (blue) ~0.72 (red)
  • seeing measured on images ~0.83
  • seeing measured on last images ~1.2 (blue) ~1 (red)

02:26 Slew to M31_AndXXVII_400s_12 * dofpia

02:30 M31_AndXXVII_400s_12
  • seeing measured on first images ~0.94 (blue) <0.94 (red)
  • seeing measured on first images <1.1
  • seeing measured on last images <0.83

02:55 Slew to M31_AndXXV_400s_13 * dofpia

03:00 M31_AndXXV_400s_13
  • seeing measured on first images <0.79
  • seeing measured on first images <0.81 (blue) <0.74 (red)
  • seeing measured on last images <0.83 (blue) <0.79 (red)

03:23 Slew to M31_AndXXVII_400s_13 * dofpia

03:29 M31_AndXXVII_400s_13
  • seeing measured on first images ~0.72 (blue) <0.79 (red)
  • seeing measured on images <0.81 (blue) ~0.77 (red)
  • seeing measured on last images ~0.77 (blue) ~0.81 (red)

03:52 Slew to M31_AndXXV_400s_14 * dofpia

03:57 M31_AndXXV_400s_14
  • seeing measured on first images ~0.72 (blue) <0.79 (red)
  • seeing measured on images ~0.72 (blue) ~0.74 (red)
  • seeing measured on last images <0.86

04:20 Slew to M31_AndXXVII_400s_14 * dofpia

04:24 M31_AndXXVII_400s_14
  • seeing measured on images < 0.87 (blue) < 0.9 (red)

04:47 Slew to M31_AndXXV_400s_15 * dofpia

04:51 M31_AndXXV_400s_15
  • seeing measured on first images ~0.7 (blue) ~0.83 (red)
  • seeing measured on images <0.81 (blue) ~0.9 (red)
  • seeing measured on last images ~0.7 (blue) 0.8-0.9 (red)

05:14 Slew to M31_AndXXVII_400s_15
  • dofpia

05:18 M31_AndXXVII_400s_15
  • seeing measured on first images ~0.74 (blue) <0.79 (red)
  • seeing measured on images ~0.74

05:41 Slew to Stetson_L92_3 for standard calibration (airmass=1.19)

05:47 Slew to Stetson_L92_1 for standard calibration (airmass=1.18)
  • OB has been modified
  • dither in 4 positions

06:02 Slew to M31_AndXXV_400s_16
  • dofpia

06:08 M31_AndXXV_400s_16
  • seeing measured on images ~1.1
  • seeing measured on last images ~ 0.97 (blue) <1.1 (red)

06:31 Slew to M31_AndXXVII_400s_16
  • dofpia

06:35 M31_AndXXVII_400s_16
  • seeing measured on first images <1 (blue) <1.1 (red)
  • seeing measured on images ~0.77 (blue) <0.79 (red)
  • seeing measured on last images <1

06:58 Slew to M31_AndXXV_400s_17
  • dofpia

07:02 M31_AndXXV_400s_17
  • seeing measured on first images <0.99 (blue) <0.97 (red)
  • seeing measured on images <0.92 (blue) <0.86 (red)
  • seeing measured on last images ~1.7 (blue)<1.4 (red)

07:25 Slew to M31_AndXXVII_400s_17
  • dofpia

07:28 M31_AndXXVII_400s_17
  • seeing measured on first images <1.58 (blue) <1.47 (red)
  • seeing measured on images <1.17 (blue) ~1.04 (red)
  • seeing measured on last images ~1.3 (blue) ~1.08 (red)

07:50 Slew to M31_AndXXV_400s_18
  • dofpia
  • pupils not found because of clouds - dofpia again (07:44)
  • converged, but it did not look good - dofpia again

08:04 M31_AndXXV_400s_18
  • seeing measured on first images ~ 1.8 (blue) ~1.6 (red)
  • seeing measured on images ~1.55 (blue) 1.4 (red)
  • seeing measured on last images <1.06 (blue) ~0.9 (red)

08:27 Slew to ngc1817
  • dofpia

08:35 ngc1817.ob
  • there were some clouds earlier and in different regions of the sky
  • seeing measured on images <0.86 (blue) <0.9 (red)
  • check images (maybe OB out of specs)

We do not repeat standard stars because the sky is not clear

08:50 SWITCH TO MODS

09:14 preset failed durinfg pointing correction

09:24 send acqMODS mods.7190.201130b1.acq
  • guider seeing ~1''
  • center the mask

09:26 while we were centering the mask the preset failed because we lost guiding

09:35 preset again - center the mask
  • the shell ripped - preset canceled

10:04 AO is up and running - repeat acquisition
  • 30min lost to AO secondary problem

10:18 execMODS mods.7190.201130b1.obs

11:13 preset failed because of rotator fault
  • only 1 of 4 exposures completed
  • 30 min lost because of LDG rotator fault

11:43 repeat preset and center mask

11:57 execMODS mods.7190.201130b1.obs

12:37 abort during the second exposure because of twilight.
  • 2 of 4 exposures completed for this mask
  • mods1r.20111022.0004.fits - 2250sec
  • mods1r.20111022.0008.fits - 2250sec

12:38 execMODS standard_star.acq/obs
  • mods1r.20111022.0011. - 0012.fits (12sec - 60sec)

12:55 Close telescope - start MODS calibrations
  • gmosflatsLBTb2.cal
  • grlampsb2.cal
  • grlampsb1.cal
  • gmosflatsLBTb1.cal
  • standard_grlamps.cal
  • grslitflatLBT.cal
  • grslitLESSflatLBT.cal
  • allbias.cal

-- KonstantinaBoutsia - 22 Oct 2011
Topic revision: r10 - 26 Nov 2011, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback